OpenStudy (errinss):

I dont like word problems>-< At 3:00 p.m., Jamie leaves her house and drives north to the beach. She arrives at 5:30 p.m. Also leaving at 3:00 p.m., her brother Raul leaves her house and drives south to a state park. He arrives at 4:30 p.m. The driving distance from the beach to the state park is 240 mi. Jamie and Raul drove at the same speed. What was their speed?

OpenStudy (errinss):

@steve816

OpenStudy (steve816):

Uh, I hate word problems too :(

OpenStudy (narissa):

I'm horrible at math! but somehow good at word problems...So Jamie took 2 hr and 30 min to reach the because 5:30-3:00=2hrs and 30 min. Raul took 1 hr 30 min to reach state park because 4:30-3:00=1 hr and 30 min.Since they went in opposite directions you can add the time to get 4 hour.It would take 4 hours to go from the beach to state park.There speed was 60 miles per hour since 240/60=4/1 mph.